nolegirl24:

Which of the following best defines 4 to the power of 2 over 3? Square root of 16 Cube root of 4 Square root of 4 Cube root of 16

Magic:

The square root of 4 is 2, and \[4^2/3=2.5\] which is the closest.

Magic:

@Vocaloid I am probably wrong, so check my answers when you get on.

nolegirl24:

Thanks for your help @Magic

nolegirl24:

How do you figure out 4 to the power of 2 over 3 though?

Magic:

4 to the power of 2 is 4^2. Then put it over 3, 4^2/3

nolegirl24:

So you do 4*4 and then divide it by 3?!?!?

Magic:

That would be 5.3. As I said, wait for Vocaloid to explain stuff. I hate math. She should be on soon.

nolegirl24:

I go the answer..... \[4^2/3 = \sqrt[3]{4^2} = \sqrt[3]{16}\]

nolegirl24:

So it would be the cube root of 16